Heat is supplied to an engine at the rate of 10^(6)J minute ^(-1) and the engine has an output of 10 horsepower.What is the efficiency of the engine and what is the heat output per minute?


1
Expert's answer
2021-04-13T06:48:58-0400

Answer

Heat supplied

Q"_1=10^6J\/min"

W=10"\times 746\times60=447600J/min"

So effeciency

"\eta=\frac{W}{Q_1}=\frac{447600}{10^6}=45" %


heat output per minute

"Q_2=Q_1-W=1000000-447600=552400J\/min"